### v2

@@ 5.2 What Works The Best?
5.3 Practical Considerations
@@
COCO 0.88 - 4.07 0.66 - 0.66 1.55 - 1.55
PPD 8.46 18.01 - 0 7.85 - 0 12.13 -
5.4 Individual Factors
In the next series of experiments, we study the effect of various elements that could influence the performance of the models. Namely, these include the model size, the underlying architecture, the pretrained dataset and the pretraining method. By studying these individual components, we seek to provide a more granular understanding of what drives model performance.
5.4.1 Input Resolution.
@@ 5.4 Individual Factors
5.4.2 Model Size.
Table 4 demonstrates that a larger model size does not translate to a big boost of performance. Specifically, for SAM and DINO V2 with the linear method, the base model (ViT-B) yields, on average, competitive performance, despite having 3 times less parameters than the (ViT-L) models. In the case of ResNet, moving from ResNet-50 to ResNet-101 shows a performance boost, though the increase is modest. For the LoRA method, fine-tuning of a subset of parameters of the encoder yields better performance with bigger models. These observations partially accounts for the performance discrepancy between SAM, DINO V2, and ResNet50, highlighting the influence of model size differences.
Table 4: Effect of model size on mIoU.
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
SAM (ViT-B) 35.72$\pm$01.50 03.13$\pm$00.09 83.14$\pm$10.36 38.50$\pm$00.12 05.85$\pm$00.13 84.94$\pm$10.11
5.4.3 Training Dataset.
To examine the impact of the training dataset on model performance, we compared DINO V1 and DINO V2, noting that the primary distinction between them lies in their training datasets. DINO V1 was trained on ImageNet-1K, whereas DINO V2 utilized a much larger dataset comprising 142M curated and uncurated images. It is important to mention that another difference between the two models is the employed loss function, although the focus of this analysis is primarily on the effect of dataset size and composition. Table 5 illustrates that DINO V2 significantly surpasses DINO V1 in performance, with DINO V1 exhibiting notably weak results on the COCO dataset. Showing that having a bigger training dataset is beneficial for the downstream task performance on few-shot semantic segmentation.
Table 5: Effect of the training dataset on mIoU
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
DINO V1 29.03$\pm$01.08 05.92$\pm$00.39 82.14$\pm$10.04 32.53$\pm$00.97 06.89$\pm$00.44 82.99$\pm$09.35
5.4.4 Architecture.
To assess the influence of model architecture, we analyzed the performance differences between DINO V1 (ViT-S) and its ResNet-50 variant, alongside comparing the CLIP (ViT-B) version against the CLIP ResNet-50 version. It is noteworthy that for CLIP, the ViT-B model is larger than its ResNet counterpart, an aspect previously discussed in terms of model size impact. According to Table 6, ViT architectures demonstrate a clear performance advantage over their ResNet equivalents. This trend underscores the potential superiority of ViT models in adapting to various tasks.
Table 6: Effect of model architecture on mIoU
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
DINO V1 (ViT-S) 26.61$\pm$00.53 04.86$\pm$00.23 81.86$\pm$08.19 29.90$\pm$00.95 05.82$\pm$00.34 79.70$\pm$09.03
5.4.5 Training Method.
To investigate the effect of different self-supervised learning (SSL) methods on model performance, we conducted a comparison between DINO V1 and MAE. Both models were trained on the same dataset (ImageNet-1K), and share identical architecture and size (ViT-B with 86M parameters), making the SSL method the primary variable. The comparison in Table 7 reveals that DINO V1 surpasses MAE on the Cityscapes and PPD datasets, whereas MAE demonstrates superior performance on COCO. This mixed outcome indicates that there is no definitive answer as to which SSL training approach offers the best adaptability for few-shot segmentation tasks.
Table 7: Effect of the training method on mIoU
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
DINO V1 29.03$\pm$01.08 05.92$\pm$00.39 82.14$\pm$10.04 32.53$\pm$00.97 06.89$\pm$00.44 82.99$\pm$09.35
MAE 23.63$\pm$00.89 08.54$\pm$00.46 74.34$\pm$10.88 26.10$\pm$00.26 10.13$\pm$00.42 72.87$\pm$12.83
5.4.6 Registers.
In the study by [5], it is demonstrated that DINO V2, unlike its predecessor DINO V1, presents artifacts within its feature maps. This phenomenon is attributed to the presence of tokens with unusually high norms, which are integral to the model’s internal computations. To address this issue, the authors introduced a modified version of DINO V2, incorporating additional tokens termed “Registers” during the training phase, which are subsequently omitted during inference. As evidenced in Table 8, this innovative approach significantly enhances DINO V2’s performance on the COCO dataset, showcasing the potential benefits of Registers during training for few-shot segmentation tasks.

### v3

@@ 6.2 What Works The Best?
SegFormer 59.53$\pm$00.52 64.08$\pm$00.51 60.99$\pm$00.70 60.21$\pm$00.29 62.46$\pm$00.56
6.3 Individual Factors
In the next series of experiments, we study the effect of various elements that could influence the performance of the models. Namely, these include the model size, the underlying architecture, the pretrained dataset, the pretraining method and the adaptation method. To isolate the effects of each component, we compare foundation models that differ primarily in one of these aspects, although minor variations may exist between the models. By studying these individual components, we seek to provide a more granular understanding of what drives model performance. (Other individual factors such as input resolution and registers are discussed in the Appendix A.5).
6.3.1 Model Size.
Table 2 demonstrates that a larger model size does not necessarily translate to a big boost of performance. Specifically, for SAM and DINOv2 with the linear method, the base model (ViT-B) yields, on average, competitive performance, despite having 3 times less parameters than the large (ViT-L) models. In the case of ResNet, moving from ResNet-50 (23M parameters) to ResNet-101 (44M parameters) shows a performance boost, though the increase is modest.
For the LoRA method, fine-tuning of a subset of parameters of the encoder yields better performance with bigger models. These observations partially accounts for the performance discrepancy between SAM, DINOv2, and ResNet50, highlighting the influence of model size differences.
For tiny models like SegFormer-B0, which has 3.7M parameters, there is a noticeable drop in performance, indicating that models adapted for device deployment struggle with poor adaptation performance. Enhancing the adaptation of these models remains an area for future research.
Table 2: Effect of different components on mIoU. Bold numbers indicate the best performance among the compared models for a specific component.
Linear LoRA
Cityscapes COCO PPD Cityscapes COCO PPD
[…]
SegFormer-B0 30.97$\pm$01.88 08.48$\pm$00.34 72.35$\pm$07.02 32.15$\pm$00.94 09.70$\pm$00.41 87.65$\pm$01.35
SegFormer-B5 40.55$\pm$03.40 18.37$\pm$01.13 77.88$\pm$07.82 42.34$\pm$02.20 19.64$\pm$00.49 87.82$\pm$03.78
[…]
CLIP 23.83$\pm$01.51 11.32$\pm$00.31 72.49$\pm$10.78 27.95$\pm$01.62 13.62$\pm$00.78 73.13$\pm$11.70
OpenCLIP 21.01$\pm$02.28 11.64$\pm$00.40 71.85$\pm$09.85 22.73$\pm$01.58 14.34$\pm$00.06 71.55$\pm$09.42
MAE (IN1k) 23.63$\pm$00.89 08.54$\pm$00.46 74.34$\pm$10.88 26.10$\pm$00.26 10.13$\pm$00.42 72.87$\pm$12.83
MAE (IG-3B) 21.73$\pm$01.07 05.76$\pm$00.16 75.18$\pm$09.26 23.59$\pm$00.21 07.84$\pm$00.37 80.05$\pm$07.44
[…]
OpenCLIP(ViT-B) 21.01$\pm$02.28 11.64$\pm$00.40 71.85$\pm$09.85 22.73$\pm$01.58 14.34$\pm$00.06 71.55$\pm$09.42
OpenCLIP(ConvNext-L) 29.29$\pm$2.80 11.30$\pm$00.02 81.05$\pm$08.30 38.52$\pm$01.14 17.05$\pm$00.20 84.94$\pm$09.72
[…]
MAE (MIM) 23.63$\pm$00.89 08.54$\pm$00.46 74.34$\pm$10.88 26.10$\pm$00.26 10.13$\pm$00.42 72.87$\pm$12.83
iBOT (KD + MIM) 29.46$\pm$01.72 09.31$\pm$00.68 85.06$\pm$08.80 32.54$\pm$01.43 11.27$\pm$00.99 85.48$\pm$07.80
6.3.2 Training Dataset.
To evaluate the impact of training dataset characteristics on model performance, we conducted a comparative analysis of DINOv1 and DINOv2. DINOv1 was trained on the standard ImageNet-1K dataset, while DINOv2 utilized a more extensive dataset of 142 million curated and uncurated images. Additionally, it is noteworthy that the models differ in their loss functions; however, our primary focus remains on the influence of dataset size and composition.
As illustrated in Table 2, DINOv2 significantly outperforms DINOv1, particularly demonstrating superior results on the COCO dataset. This suggests that a larger training dataset may enhance performance in few-shot semantic segmentation tasks. Nonetheless, this performance boost could also be influenced by other factors differentiating DINOv1 from DINOv2.
Furthermore, we compared CLIP, trained on 400 million (image, text) pairs, with OpenCLIP (Cherti et al., 2023) trained on the LAION-5B (Schuhmann et al., 2022) dataset comprising 5 billion (image, text) pairs, and MAE trained on ImageNet-1K against another MAE variant trained on the Instagram-3B dataset (Singh et al., 2023) , which includes 3 billion images. According to Table 2, for MAE and CLIP, an increase in training dataset size does not consistently result in better performance and can sometimes lead to a performance decline on certain datasets. Overall, the benefits of larger pre-training datasets remain ambiguous, indicating that the advantages of scaling up the training data are not universally guaranteed.
6.3.3 Architecture.
To assess the influence of model architecture, we analyzed the performance differences between DINOv1 (ViT-S) and its ResNet-50 variant, alongside comparing the CLIP (ViT-B) version against the CLIP ResNet-50 version. It is noteworthy that for CLIP, the ViT-B model is larger than its ResNet counterpart, an aspect previously discussed in terms of model size impact. According to Table 2, ViT architectures demonstrate a clear performance advantage over their ResNet equivalents. This trend underscores the potential superiority of ViT models in adapting to various tasks. For OpenCLIP, the ConvNext-L outperforms the ViT-B, this can be explained by the fact that we used a resolution of 1120 for ConvNext as in (Yu et al., 2023), and that CLIP only uses a resolution of 224 as discussed in A.5.
6.3.4 Training Method.
To investigate the effect of different self-supervised learning (SSL) methods on model performance, we conducted a comparison between knowledge distillation (KD) and masked image modeling (MIM) models. Despite our initial expectations that MIM models would surpass KD models (Park et al., 2023), the comparison in Table 2 reveals that DINOv1 (KD) surpasses MAE (MIM) on the Cityscapes and PPD datasets, and MAE outperforms DINOv1 on COCO. This mixed outcome indicates that there is no definitive answer as to which SSL training approach offers the best adaptability for FSS tasks. However, iBOT, which integrates both KD and MIM techniques, shows superior performance over both models, leveraging the strengths of both approaches.
7 Conclusion
We have introduced a novel benchmark specifically designed to assess the adaptation of Vision Foundation Models (VFM) to Few-Shot Semantic Segmentation (FSS). Our evaluation of multiple VFMs across this new benchmark has highlighted several key areas for further exploration : DINOv2 emerges as the top-performing backbone model, consistently outperforming other foundation models. While SAM exhibits strong overall results, it presents unexpected limitations on certain datasets. Surprisingly, our experiments underscore the competitive advantages of straightforward and efficient methods such as simple linear segmentation heads, challenging the necessity of complex procedures?
This work not only presents a comparison of techniques but also a more realistic and practical benchmark. We hope this benchmark will enhance the development of few-shot semantic segmentation methods and offer a robust framework for future comparisons.
Appendix A Appendix
A.1 Detailed Results for Every Dataset
We can see in Tables 3, 4, 5, 6 the detailed results on the 3 datasets separately:
  • The range of performance on the 3 datasets is different mainly due to the varying size and number of classes. On PPD, we only have 2 classes making the task easier but the training is more unstable due to the restricted number of training examples, since we only have 2 images in our support set. This leads to a bigger standard deviation in PPD compared to COCO and Cityscapes since the choice of these 2 images is crucial and the images are different between the 3 folders constituting the dataset.
  • SAM image encoder performance is poor on COCO compared to other models even though it yielded good performance in the other 2 datasets. This is probably due to the fact that we isolated the image encoder (the adaptation of the whole SAM model to prompt free Few-Shot Semantic Segmentation is left to future work).
Table 3: Detailed results of mIoU on our three considered datasets in 1-shot semantic segmentation.
Linear Multilayer SVF LoRA Fine-tuning
[…]
DINOv2 48.80$\pm$01.82 46.77$\pm$02.80 51.96$\pm$01.37 54.35$\pm$01.55 50.87$\pm$01.25
iBOT 29.46$\pm$01.72 26.84$\pm$02.04 32.16$\pm$01.20 32.54$\pm$01.43 31.78$\pm$01.31
CLIP 23.83$\pm$01.51 26.10$\pm$02.01 28.72$\pm$01.06 27.95$\pm$01.62 27.74$\pm$01.38
MAE 23.63$\pm$00.89 23.81$\pm$00.76 25.88$\pm$00.32 26.10$\pm$00.26 25.07$\pm$00.64
FCN-ResNet50 32.39$\pm$01.79 34.12$\pm$02.01 36.14$\pm$01.19 35.74$\pm$01.41 36.17$\pm$01.41
SegFormer 40.55$\pm$03.40 44.22$\pm$03.50 42.52$\pm$01.97 42.34$\pm$02.20 42.19$\pm$02.63
[…]
DINOv2 23.24$\pm$00.38 20.92$\pm$00.19 28.30$\pm$00.77 28.99$\pm$01.33 28.15$\pm$01.24
iBOT 09.31$\pm$00.68 09.35$\pm$00.47 10.91$\pm$01.02 11.27$\pm$00.99 11.67$\pm$01.09
CLIP 11.32$\pm$00.31 14.47$\pm$00.29 15.70$\pm$00.71 13.62$\pm$00.78 15.13$\pm$00.91
MAE 8.54$\pm$00.46 08.60$\pm$00.45 10.52$\pm$00.45 10.13$\pm$00.42 10.71$\pm$00.57
FCN-ResNet50 10.08$\pm$00.32 08.82$\pm$00.88 12.05$\pm$00.33 11.71$\pm$00.24 11.88$\pm$00.38
SegFormer 18.37$\pm$01.13 19.69$\pm$00.82 21.25$\pm$00.52 19.64$\pm$00.49 21.16$\pm$00.64
[…]
DINOv2 92.31$\pm$01.84 92.85$\pm$01.76 93.05$\pm$02.57 89.67$\pm$06.43 92.61$\pm$02.10
iBOT 85.06$\pm$08.80 87.52$\pm$06.23 86.04$\pm$09.37 85.48$\pm$07.80 87.63$\pm$08.00
CLIP 72.49$\pm$10.78 78.14$\pm$06.81 78.28$\pm$12.63 73.13$\pm$11.70 73.19$\pm$08.69
MAE 74.34$\pm$10.88 77.54$\pm$07.95 76.90$\pm$07.95 72.87$\pm$12.83 73.62$\pm$10.59
FCN-ResNet50 70.86$\pm$05.91 89.31$\pm$01.88 82.01$\pm$08.04 76.91$\pm$10.29 70.84$\pm$06.15
SegFormer 77.88$\pm$07.82 82.98$\pm$07.00 88.99$\pm$05.98 87.82$\pm$03.78 90.14$\pm$03.72
Table 4: Detailed results of mIoU on our three considered datasets in 2-shot semantic segmentation.
Linear Multilayer SVF LoRA Fine-tuning
[…]
DINOv2 56.54$\pm$00.75 53.36$\pm$00.09 58.21$\pm$01.00 58.84$\pm$01.75 58.60$\pm$00.57
iBOT 33.25$\pm$00.62 32.42$\pm$00.58 35.60$\pm$00.85 36.37$\pm$00.93 35.29$\pm$00.87
CLIP 25.19$\pm$01.05 28.96$\pm$01.62 30.85$\pm$00.42 30.53$\pm$01.19 19.29$\pm$04.90
MAE 24.15$\pm$00.32 24.88$\pm$00.51 27.68$\pm$00.21 27.92$\pm$00.51 26.83$\pm$00.32
FCN-ResNet50 36.40$\pm$00.57 37.89$\pm$00.45 38.98$\pm$00.06 38.34$\pm$00.28 38.97$\pm$00.33
SegFormer 45.42$\pm$00.82 50.68$\pm$01.12 46.88$\pm$01.35 47.35$\pm$01.34 46.96$\pm$01.72
[…]
DINOv2 33.08$\pm$00.17 30.01$\pm$00.57 39.91$\pm$00.90 39.01$\pm$01.36 40.09$\pm$00.68
iBOT 13.13$\pm$00.81 13.13$\pm$00.63 15.65$\pm$00.75 16.84$\pm$00.78 16.84$\pm$00.88
CLIP 16.03$\pm$00.47 18.85$\pm$00.67 22.33$\pm$00.46 20.40$\pm$00.99 21.85$\pm$00.84
MAE 12.20$\pm$00.69 12.26$\pm$00.66 15.30$\pm$00.27 15.26$\pm$00.34 15.66$\pm$00.21
FCN-ResNet50 12.54$\pm$00.50 11.44$\pm$00.48 13.23$\pm$00.78 13.42$\pm$00.81 12.55$\pm$00.83
SegFormer 24.23$\pm$00.20 25.38$\pm$00.25 26.34$\pm$01.34 25.62$\pm$00.75 28.08$\pm$00.72
[…]
DINOv2 93.96$\pm$01.44 94.24$\pm$02.03 94.17$\pm$01.68 93.33$\pm$02.18 94.35$\pm$00.86
iBOT 90.43$\pm$05.44 93.05$\pm$02.63 90.79$\pm$06.09 90.23$\pm$06.42 91.59$\pm$04.22
CLIP 76.01$\pm$10.95 82.58$\pm$03.02 83.59$\pm$05.94 77.21$\pm$09.11 65.92$\pm$15.61
MAE 74.36$\pm$12.50 79.05$\pm$07.61 78.12$\pm$11.30 77.93$\pm$12.92 74.97$\pm$10.59
FCN-ResNet50 74.01$\pm$05.93 92.74$\pm$02.02 93.76$\pm$02.01 87.88$\pm$06.22 94.91$\pm$00.59
SegFormer 82.74$\pm$04.82 90.67$\pm$06.62 90.64$\pm$04.80 85.66$\pm$02.75 91.39$\pm$01.59
Table 5: Detailed results of mIoU on our three considered datasets in 5-shot semantic segmentation.
Linear Multilayer SVF LoRA Fine-tuning
[…]
DINOv2 59.71$\pm$02.30 59.22$\pm$02.16 61.22$\pm$01.97 61.69$\pm$01.37 62.88$\pm$01.74
iBOT 36.45$\pm$00.24 36.67$\pm$00.22 39.48$\pm$00.24 41.05$\pm$00.72 40.46$\pm$00.47
CLIP 28.70$\pm$01.20 31.21$\pm$01.62 33.88$\pm$01.23 33.23$\pm$00.78 13.40$\pm$01.31
MAE 27.49$\pm$00.72 28.19$\pm$01.01 30.67$\pm$00.91 30.92$\pm$00.96 29.91$\pm$01.03
FCN-ResNet50 37.68$\pm$00.97 39.85$\pm$00.13 42.42$\pm$00.17 42.80$\pm$00.43 43.30$\pm$00.75
SegFormer 48.26$\pm$01.13 53.39$\pm$01.32 48.00$\pm$00.75 50.06$\pm$00.49 50.00$\pm$00.63
[…]
DINOv2 44.48$\pm$00.64 42.32$\pm$00.66 49.07$\pm$01.59 45.82$\pm$01.80 50.52$\pm$01.34
iBOT 20.99$\pm$00.81 20.85$\pm$00.61 24.33$\pm$00.96 27.62$\pm$00.88 27.28$\pm$01.20
CLIP 26.71$\pm$00.57 29.11$\pm$00.67 31.85$\pm$00.15 30.71$\pm$00.11 31.61$\pm$00.39
MAE 19.20$\pm$00.52 19.27$\pm$00.44 21.67$\pm$00.46 22.80$\pm$00.40 23.39$\pm$00.39
FCN-ResNet50 15.42$\pm$00.27 14.23$\pm$00.65 15.47$\pm$00.23 15.94$\pm$00.16 14.09$\pm$00.10
SegFormer 31.80$\pm$01.20 33.47$\pm$01.19 34.68$\pm$01.47 32.87$\pm$01.23 36.56$\pm$01.52
[…]
DINOv2 95.21$\pm$00.35 96.13$\pm$00.35 95.62$\pm$00.47 95.11$\pm$00.39 93.06$\pm$03.70
iBOT 95.02$\pm$00.95 95.71$\pm$00.25 95.54$\pm$00.63 95.34$\pm$00.71 96.57$\pm$00.16
CLIP 83.48$\pm$01.12 85.68$\pm$00.43 88.94$\pm$00.58 83.07$\pm$01.68 85.76$\pm$01.16
MAE 86.39$\pm$00.42 87.61$\pm$00.31 86.26$\pm$01.99 87.34$\pm$00.74 86.73$\pm$00.36
FCN-ResNet50 87.97$\pm$00.79 95.02$\pm$01.11 96.84$\pm$00.15 95.45$\pm$00.65 97.13$\pm$0.02
SegFormer 90.48$\pm$00.82 95.92$\pm$00.67 94.56$\pm$00.06 91.97$\pm$00.24 94.03$\pm$00.19
Table 6: Detailed results of mIoU on our three considered datasets in 10-shot semantic segmentation.
Linear Multilayer SVF LoRA Fine-tuning
[…]
DINOv2 63.19$\pm$00.23 63.64$\pm$00.64 63.38$\pm$00.45 63.87$\pm$00.46 66.24$\pm$00.49
iBOT 39.47$\pm$00.51 39.22$\pm$00.98 40.46$\pm$00.80 43.05$\pm$01.07 43.79$\pm$01.05
CLIP 32.44$\pm$01.32 34.83$\pm$00.82 36.76$\pm$01.07 36.50$\pm$01.17 21.83$\pm$10.21
MAE 30.46$\pm$00.28 30.59$\pm$00.18 32.63$\pm$00.75 32.99$\pm$00.45 31.85$\pm$00.55
FCN-ResNet50 40.63$\pm$00.66 43.20$\pm$00.38 45.12$\pm$00.53 45.56$\pm$00.62 45.63$\pm$01.38
SegFormer 50.76$\pm$00.56 56.77$\pm$00.66 50.12$\pm$00.45 52.81$\pm$00.03 52.08$\pm$00.17
[…]
DINOv2 50.84$\pm$00.43 49.83$\pm$00.34 50.84$\pm$01.23 48.55$\pm$00.66 52.74$\pm$01.09
iBOT 28.23$\pm$00.46 28.30$\pm$00.38 29.03$\pm$00.83 35.26$\pm$01.08 35.20$\pm$00.80
CLIP 33.75$\pm$00.36 36.45$\pm$00.53 35.36$\pm$00.54 35.38$\pm$00.33 34.19$\pm$00.10
MAE 25.38$\pm$00.22 25.88$\pm$00.29 20.96$\pm$00.54 08.84$\pm$11.26 26.86$\pm$00.25
FCN-ResNet50 17.04$\pm$00.43 16.74$\pm$00.46 16.81$\pm$00.24 17.15$\pm$00.21 15.55$\pm$00.17
SegFormer 36.47$\pm$00.91 38.97$\pm$00.88 37.64$\pm$01.66 35.59$\pm$00.56 40.61$\pm$01.32
[…]
DINOv2 95.55$\pm$00.15 96.38$\pm$00.24 95.67$\pm$00.53 95.32$\pm$00.41 94.24$\pm$02.64
iBOT 95.53$\pm$00.08 96.02$\pm$00.13 96.01$\pm$00.11 95.68$\pm$00.22 96.78$\pm$00.11
CLIP 84.24$\pm$00.51 86.18$\pm$00.27 88.81$\pm$00.76 83.37$\pm$00.74 86.84$\pm$00.39
MAE 86.84$\pm$00.74 87.44$\pm$00.48 86.59$\pm$00.46 87.00$\pm$01.90 87.17$\pm$00.45
FCN-ResNet50 88.40$\pm$01.72 96.01$\pm$00.28 97.18$\pm$00.20 96.34$\pm$00.16 97.39$\pm$00.11
SegFormer 91.35$\pm$00.13 96.49$\pm$00.06 95.21$\pm$00.15 92.21$\pm$00.30 94.68$\pm$00.27
A.2 Parameter Efficient Fine-Tuning (PEFT) Methods.
Given the strong performance of Parameter-Efficient Fine-Tuning (PEFT) methods on our benchmark, we opted to evaluate various PEFT approaches from multiple families. We examined LoRA and SVF alongside two adapter tuning methods: AdaptFormer (Chen et al., 2022), which integrates a lightweight adapter module parallel to the feed-forward network within the encoder blocks and only trains these adapters, and Convpass (Jie & Deng, 2022), which adds convolutional layers to the parallel adapter to incorporate inductive biases and trains only these layers. Additionally, we assessed Visual Prompt Tuning (VPT) (Jia et al., 2022), which introduces tokens at the input of encoder blocks and trains only these tokens, and a partial-based tuning method, BitFit(Ben Zaken et al., 2022), which freezes the model weights and fine-tunes only the biases. For computing constraints, we only tested these methods with DINOv2 on 1-shot setting.
The results, presented in Table 7, indicate minimal differences among the PEFT methods, suggesting no definitive superior method. This outcome underscores the critical role of feature extractors in determining the effectiveness of these tuning approaches.
Table 7: Effect of PEFT methods on mIoU, bold numbers denote top performance across methods.
SVF LoRA VPT Adaptformer Convpass Bitfit
Cityscapes 51.96$\pm$01.37 54.35$\pm$01.55 51.92$\pm$00.56 50.44$\pm$00.53 50.24$\pm$01.12 51.55$\pm$00.58
COCO 28.30$\pm$00.77 28.99$\pm$01.33 30.15$\pm$02.10 28.68$\pm$01.23 29.35$\pm$01.63 29.19$\pm$00.94
PPD 93.05$\pm$02.57 89.67$\pm$06.43 91.25$\pm$02.11 92.87$\pm$01.47 92.21$\pm$02.50 92.91$\pm$01.50
A.3 SAM Underperformance
  • Decoder Limitations: The primary suspect for SAM’s subpar performance is the dependency of the image encoder on the prompt encoder and the mask decoder. To address this, we replaced the original mask decoder with a transformer decoder, as detailed in (Kirillov et al., 2023), modifying it for few-shot semantic segmentation following the inspiration from (Zhong et al., 2024). We introduced a constant prompt to the decoder, altering tokens to produce an $N$-class depth map rather than binary masks, and eliminated the IoU score output branch. We experimented with two mask decoder configurations, one with a Conv Transpose module and another where we upscaled the output through interpolation, while initializing remaining modules with pretrained weights. Despite these adjustments, this approach was outperformed by a simpler linear decoding method by up to 5% mIoU, likely due to the challenges of training a complex decoder in a few-shot scenario.
  • Mask Distribution Bias: Another investigated issue was the mask distribution bias in COCO towards image centers, unlike the distribution in the SA-1B dataset used for SAM, highlighted in (Kirillov et al., 2023), where masks are more evenly distributed. We attempted to mitigate this bias with various preprocessing techniques (RandomCrop, CenterCrop, dividing each image into four smaller images); however, these modifications did not enhance our results.
A.4 Practical Considerations
@@
COCO 00.88 - 04.07 00.66 - 00.66 01.55 - 01.55
PPD 08.46 18.01 - 0 07.85 - 0 12.13 -
[…]
COCO 0 - 00.35 0 - 00.84 00.65 - 0
PPD 01.97 01.97 - 01.67 01.67 - 02.01 0 -
[…]
COCO 0 - 0 0 - 0 0 - 0
PPD 0 0 - 0 0 - 0 0 -
A.5 More Individual Factors
A.5.1 Input Resolution Effect.
@@ A.5 More Individual Factors
A.5.2 Registers Effect.
In the study by (Darcet et al., 2023), it is demonstrated that DINOv2, unlike its predecessor DINOv1, presents artifacts within its feature maps. This phenomenon is attributed to the presence of tokens with unusually high norms, which are integral to the model’s internal computations. To address this issue, the authors introduced a modified version of DINOv2, incorporating additional tokens termed “Registers” during the training phase, which are subsequently omitted during inference. As evidenced in Table 9, this innovative approach significantly enhances DINOv2’s performance on the COCO dataset, showcasing the potential benefits of Registers during training for FSS tasks.